OpenStudy (anonymous):

Which of the following is equivalent to log(16)8?

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

anyone?

Nnesha (nnesha):

which of the following? i guess you have to apply change of base formula which is \[\huge\rm log_a b = \frac{ \log b }{ \log a }\]

OpenStudy (anonymous):

is this log 8 to base 16? or log 16 to base 8? what are the multiply choices?

OpenStudy (anonymous):

a 2 b.1.248 c. 4/3 d. 3/4

Nnesha (nnesha):

i guess its \[\huge\rm log_{16} 8\] bec 16 in parentheses

Nnesha (nnesha):

apply change of base formula

OpenStudy (lυἶცἶ0210):

You have to rewrite this in exponential form \(\LARGE \log_{a}b=c \) is the same as \(\LARGE a^c=b \) so... \(\LARGE \log_{16} 8=x\) is the same as \(\LARGE 16^x=8 \)

OpenStudy (kitkat16):

0.75 = 3/4